OpenStudy (anonymous):

Explain how applying the Associative Property of Addition to 7.5+(2.5+4.9) makes it easier to simplify. Show the solution in your explanation.

hartnn (hartnn):

Associative property is :a+(b+c) = (a+b) + c so u can write 7.5+(2.5+4.9) as (7.5+2.5) +4.9 Now, 7.5+2.5 is easier to calculate (=10) than 2.5+4.9 hence associative property makes it easier to simplify.